if the side of a square is equal to 3,what is the area of the square?Suppose the area is givern as 16sq.m.What is the measure of its side?How do you get the answer?
Answer by stanbon(75887)

if the side of a square is equal to 3,what is the area of the square?
---
Area of a square = (side)^2
A = 3^2
A = 9 sq. units
---------------------------
Suppose the area is given as 16 sq.m.
What is the measure of its side?
How do you get the answer?
---
A = s^2
Solve 16 = s^2
---
Take the square root of both sides to get:
s = sqrt(16)
side = 4 m
===============
Cheers,
Stan H.
----